Grade 12th passAlgebra

The sum of all real values of x satisfying the equation(x^2-5x+5)^x^2+4x-60=1

Shoaib akhter
7 Years agoGrade 12th pass

1 Answer

Apoorv Tiwari
7 Years ago
as, (x^2-5x+5)^x^2+4x-60 = 1 = (x^2-5x+5)^0,
             => (x^2+4x-60) = 0
             => x = -10 and 6
and, 
(x^2-5x+5) = 0 or -1
 
now if (x^2-5x+5) = 0
then it wont satisfy original eqn,
 
if (x^2-5x+5) = 1
x = 4 and 1
if (x^2-5x+5) = -1
x = 2 and 3
 
now x =3 doesn't satisfy the equation
 
Hence solutions are -10,6,4,1 and 2
so the sum of solutions = -10+6+4+1+2 = 3
 
Answer: 3